# Preference Optimization by Estimating the Ratio of the Data Distribution

### v1

@@ Abstract
Direct preference optimization (DPO) is widely used as a simple and stable method for aligning large language models (LLMs) with human preferences. This paper investigates a generalized DPO loss that enables a policy model to match the target policy from a likelihood ratio estimation perspective. The ratio of the target policy provides a unique identification of the policy distribution without relying on reward models or partition functions. This allows the generalized loss to retain both simplicity and theoretical guarantees, which prior work such as $f$-PO fails to achieve simultaneously. We propose Bregman preference optimization (BPO), a generalized framework for ratio matching that provides a family of objective functions achieving target policy optimality. BPO subsumes DPO as a special case and offers tractable forms for all instances, allowing implementation with a few lines of code. We further develop scaled Basu’s power divergence (SBA), a gradient scaling method that can be used for BPO instances. The BPO framework complements other DPO variants and is applicable to target policies defined by these variants. In experiments, unlike other probabilistic loss extensions such as $f$-DPO or $f$-PO, which exhibit a trade-off between generation fidelity and diversity, instances of BPO improve both win rate and entropy compared with DPO. When applied to Llama-3-Instruct-8B, BPO achieves state-of-the-art performance among Llama-3-8B backbones, with a 55.9% length-controlled win rate on AlpacaEval2.
@@ 1 Introduction
DPO is an alignment method that does not require an auxiliary reward model, thereby improving training efficiency and stability. DPO training reduces to logistic regression on offline preference datasets, following the Bradley–Terry model bradley1952rank . Subsequent studies have proposed extensions of the DPO loss functions tang2024generalized ; wang2024beyond ; xu2025fpo . From the perspective of distribution matching, $f$-PO ji2024towards ; xu2025fpo reformulates the alignment objective as matching the policy model to the optimal policy defined by DPO, extending the loss using $f$-divergence. While this reformulation retains the target optimality, $f$-PO introduces additional complexity since it relies on a learned reward model and requires Monte Carlo estimation of partition functions. Importantly, its optimality cannot be guaranteed without incurring additional computational cost, as summarized in the third and fourth rows of Table 1.
@@
We empirically evaluate several instances of BPO against baselines, measuring both generation fidelity (win rate with GPT-4 judge) and generation diversity. In contrast to other probabilistic loss extensions such as $f$-DPO wang2024beyond and $f$-PO xu2025fpo , which show a clear trade-off between win rate and diversity, BPO instances improve both win rate and entropy over DPO, as summarized in Figure 1. The proposed SBA instance achieves the best trade-off between win rate and diversity among all instances. We also apply the proposed loss to Llama3-8B-Instruct grattafiori2024llama and achieve a 55.9% length-controlled win rate against GPT-4 Turbo on AlpacaEval2 alpaca_eval . To the best of our knowledge, this is state-of-the-art performance among Llama-3-8B backbone models.
Table 1: Summary of probabilistic DPO extensions. O: optimality preservation, S: simplicity without extra training cost, G: generality for multiple objectives. Refer to Appendix B for baseline details.
@@ 2.1 Alignment of large language models with human preference
Reinforcement learning from human feedback (RLHF) The alignment method RLHF bai2022training ; ouyang2022training ; stiennon2020learning ; ziegler2019fine consists of two phases following SFT. The first phase involves learning a reward model $r_{\phi}$. This phase requires access to an offline pairwise preference dataset, consisting of tuples $({\mathbf{x}}:\text{prompt},{\mathbf{y}}_{w}:\text{preferred response},{\mathbf{y}}_{l}:\text{dispreferred response})$ drawn from the underlying preference data distribution $p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}|{\mathbf{x}})p_{\text{prompt}}({\mathbf{x}})$. The learning objective of the reward model is
@@
Direct preference optimization (DPO) The alignment method DPO rafailov2023direct unifies the objectives Eqs. 1 and 2 into a single objective without the reward model. Eq. 2 has a closed-form solution:
@@
where $Z({\mathbf{x}})=\sum_{{\mathbf{y}}}\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})\text{exp}\big{(}\frac{1}{\beta}r_{\phi^{*}}({\mathbf{x}},{\mathbf{y}})\big{)}$ is the partition function. From Eq. 3, DPO defines a new reward model $r_{\text{DPO}}({\mathbf{x}},{\mathbf{y}})=\beta\log\frac{\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})}+\beta\log Z({\mathbf{x}})$, which is parameterized by the policy and reference models. Substituting $r_{\text{DPO}}$ into $\mathcal{L}_{\text{reward}}(r_{\text{DPO}};p_{\text{data}})$ yields the following objective:
@@
which enables the policy model to learn directly from a preference dataset in a supervised manner. The optimal policy $\pi_{\theta^{*}}$ that minimizes $\mathcal{L}_{\text{DPO}}$ matches the optimal solution in Eq. 3, and it satisfies:
$p_{\textrm{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}|{\mathbf{x}})=\sigma\left(\beta\log{\frac{\pi_{\theta^{*}}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\textrm{ref}}({\mathbf{y}}_{w}|{\mathbf{x}})}}-\beta\log{\frac{\pi_{\theta^{*}}({\mathbf{y}}_{l}|{\mathbf{x}})}{\pi_{\textrm{ref}}({\mathbf{y}}_{l}|{\mathbf{x}})}}\right)$ (5)
@@ 2.2 Likelihood ratio estimation under Bregman divergence
where $h$ denotes a strictly convex and twice continuously differentiable function with derivative function $h^{\prime}$, and $B_{h}$ is the pointwise Bregman divergence which measures the error of the linear approximation as illustrated in Figure 2. Specific instances are summarized in Table 2, showing that previous likelihood ratio estimation methods differ only by the choice of $h$. Among various instances, Basu’s power (BA) divergence basu1998robust , defined for $\lambda>-1$, smoothly interpolates between KLIEP (at $\lambda=0$) and LSIF (at $\lambda=1$). Inspired by the success of generalized likelihood ratio estimation in recent generative modeling studies lou2024discrete ; pmlr-v202-kim23i ; choi2021featurized , we apply this extension to generalize the DPO loss.
@@ 3 Methods
This section introduces the proposed Bregman Preference Optimization (BPO). Section 3.1 reformulates the DPO objective as a likelihood ratio estimation problem. Section 3.2 extends the DPO objective via Bregman divergence under the likelihood ratio estimation perspective. Section 3.3 analyzes instances of BPO and introduces the scaled Basu’s power (SBA) divergence within this framework. Finally, Section 3.4 discusses the applicability of BPO to other DPO variants in an orthogonal manner.
@@ 3.1 Preference optimization as likelihood ratio estimation
In contrast to prior works ji2024towards ; xu2025fpo that characterize the optimal policy using a learned reward model $r_{\phi^{*}}$ and a partition function $Z({\mathbf{x}})$ as in Eq. 3, we aim to characterize the optimal policy without such modeling complexity.
@@
Please refer to Section A.1 for the proof. Eq. 7 shows that the likelihood ratio of optimal policy $\pi_{\theta^{*}}$ can be specified solely using the reference model $\pi_{\text{ref}}$ and the preference data distribution $p_{\text{data}}$. The ratio $\frac{\pi_{\theta^{*}}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\theta^{*}}({\mathbf{y}}_{l}|{\mathbf{x}})}$ is equivalent to the concrete score meng2022concrete ; lou2024discrete (up to a constant), and the concrete score satisfies the completeness property lyu2012interpretation ; meng2022concrete . This means that the concrete score can uniquely identify the distribution $\pi_{\theta^{*}}$. Therefore, the matching $\frac{\pi_{\theta}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\theta}({\mathbf{y}}_{l}|{\mathbf{x}})}$ to $\frac{\pi_{\theta^{*}}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\theta^{*}}({\mathbf{y}}_{l}|{\mathbf{x}})}$ is sufficient for $\pi_{\theta}$ to recover $\pi_{\theta^{*}}$, providing a theoretical justification for ratio matching. To build a data-driven estimator for this matching, we rearrange Eq. 7, as detailed in Eq. 15 to Eq. 19, and reformulate preference optimization as a matching problem between data ratio $R_{\text{data}}$ and model ratio $R_{\theta}$, defined as:
@@ 3.2 Bregman preference optimization via ratio matching
As discussed in Section 3.1, preference optimization can be formulated as a ratio matching between $R_{\text{data}}$ and $R_{\theta}$. We define the loss using the Bregman ratio matching framework in Section 2.2 as:
@@
Please see Section A.2 for the proof. Theorem 2 implies that the objective $D_{h}(R_{\text{data}}||R_{\theta})$ guarantees target optimality under a valid $h$, but it is intractable since evaluating $R_{\text{data}}$ at a given point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ is infeasible. As preference optimization provides only samples from $p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}\mid{\mathbf{x}})$ without direct access to the distribution, we propose a tractable alternative inspired by implicit score matching hyvarinen2005estimation :
@@
The proof is provided in Section A.3. The intractable term $R_{\text{data}}$ does not appear in $\mathcal{L}^{h}_{\textrm{BPO}}$, and the statistical information of $R_{\text{data}}$ is captured entirely via samples from $p_{\textrm{data}}$. As the value of function $R_{\theta}$ at the point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ computed through the feed-forward passes of $\pi_{\theta}$ and $\pi_{\text{ref}}$, $R_{\theta}$ is appropriate to retain it inside the expectation. To analyze the learning dynamics of $\mathcal{L}^{h}_{\textrm{BPO}}(R_{\theta};p_{\textrm{data}})$, we provide the following gradient analysis:
@@
See Section A.4 for the proof. Eq. 11 shows that the gradient direction at a point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ depends only on $\nabla_{\theta}R_{\theta}$, regardless of the choice of $h$. In contrast, $h$ controls the point-wise gradient magnitude $G_{h}(R_{\theta})$, which determines the relative weighting of each sample during optimization. The weighting can influence the gradient direction aggregated over a mini-batch as shown in Figure 3(c). Proposition 4 further explains that gradient descent updates decrease the value of $R_{\theta}$ at the observed point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})\sim p_{\text{data}}$. This is intuitive since the observed pair is sampled from the distribution in the denominator of the target ratio $R_{\text{data}}=\frac{p_{\text{data}}({\mathbf{y}}_{w}\prec{\mathbf{y}}_{l}\mid{\mathbf{x}})}{p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}\mid{\mathbf{x}})}$. Figures 3(a) and 3(b) show that $G_{h}(R_{\theta})$ remains positive across various choices of $h$. Although Theorem 2 shows that the theoretical optimality holds regardless of the choice of the function $h$, the gradient magnitude $G_{h}(R_{\theta})$ varies with $h$, making the choice of $h$ important due to potential sub-optimality in practical optimization.
@@ 3.3 Instances of Bregman preference optimization
Table 2 summarizes well-known instances of Bregman divergences unified in sugiyama2012density and their corresponding BPO objectives $\mathcal{L}^{h}_{\text{BPO}}$. The proposed $\mathcal{L}^{h}_{\text{BPO}}$ recovers the original DPO objective $\mathcal{L}_{\text{DPO}}$ when $h$ corresponds to the logistic regression, denoted as $\mathcal{L}^{\text{LR}}_{\text{BPO}}$ (See Section A.5 for details). The instances, $\mathcal{L}^{\text{KLIEP}}_{\text{BPO}}$ and $\mathcal{L}^{\text{LSIF}}_{\text{BPO}}$, are unified under $\mathcal{L}^{\text{BA}_{\lambda}}_{\text{BPO}}$, where the gradient magnitude is given by:
@@
as shown in Figure 3(a) for various values of $\lambda$. While the term $R_{\theta}^{\lambda}+R_{\theta}^{-\lambda-1}$ provides meaningful control over the optimization behavior with respect to $R_{\theta}$ via the hyperparameter $\lambda$, the coefficient $(\lambda+1)$ unnecessarily scales up the gradient magnitude without introducing any $\theta$-dependent learning signal. Compared to the gradient magnitude of DPO, given by $G_{\text{LR}}(R_{\theta})=\frac{1}{1+R_{\theta}}$, the gradient magnitude of $G_{\text{BA}_{\lambda}}(R_{\theta})$ increases significantly as $\lambda$ becomes larger. When the gradient scale varies, training requires careful adjustment of the learning rate, batch size, and optimizer. Since preference optimization is sensitive to hyperparameters, improper tuning can severely degrade performance. To mitigate this issue, we propose a simple scaled version of BA to match DPO’s gradient scale.
Scaled Basu’s power divergence (SBA) We propose $h(R)=\frac{R^{1+\lambda}-R}{s\lambda(\lambda+1)}$, which corresponds to scaling the $h$-function of BA by a factor of $(\lambda+1)s$, where $s$ is a scaling constant. This scaling directly affects the gradient magnitude as follows:
@@
$G_{\text{SBA}_{\lambda}}(R_{\theta})$ eliminates the unnecessary $\lambda$-dependent amplification, resulting in a more reasonable gradient scale in Figure 3(b). Since preference optimization typically initializes $\pi_{\theta}$ as $\pi_{\text{ref}}$, the value of $R_{\theta}$ is 1 for all input points $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ at the start of training. By setting $s=4$, the gradient magnitude of $G_{\text{SBA}_{\lambda}}$ matches that of $G_{\text{LR}}$ at initialization ($R_{\theta}=1$). The hyperparameter $\lambda$ allows the model to control whether to prioritize updates for more confident (e.g., $R_{\theta}\ll 1$ ) or less confident (e.g., $R_{\theta}\approx 1$) samples. Recent studies wu2024beta ; kim2025spread have shown that DPO is sensitive to data quality and that applying confidence-based adjustments to individual samples can be heuristically effective. SBA controls the sensitivity to confident samples by tuning $\lambda$, while preserving theoretical optimality.

### v2

@@ Abstract
Direct preference optimization (DPO) is widely used as a simple and stable method for aligning large language models (LLMs) with human preferences. This paper investigates a generalized DPO loss that enables a policy model to match the target policy from a likelihood ratio estimation perspective. The ratio of the target policy provides a unique identification of the policy distribution without relying on reward models or partition functions. This allows the generalized loss to retain both simplicity and theoretical guarantees, which prior work such as $f$-PO fails to achieve simultaneously. We propose Bregman preference optimization (BPO), a generalized framework for ratio matching that provides a family of objective functions achieving target policy optimality. BPO subsumes DPO as a special case and offers tractable forms for all instances, allowing implementation with a few lines of code. We further develop scaled Basu’s power divergence (SBA), a gradient scaling method that can be used for BPO instances. The BPO framework complements other DPO variants and is applicable to target policies defined by these variants. In experiments, unlike other probabilistic loss extensions such as $f$-DPO or $f$-PO, which exhibit a trade-off between generation fidelity and diversity, instances of BPO improve both win rate and entropy compared with DPO. When applied to Llama-3-8B-Instruct, BPO achieves state-of-the-art performance among Llama-3-8B backbones, with a 55.9% length-controlled win rate on AlpacaEval2. Project page: https://github.com/aailab-kaist/BPO.
@@ 1 Introduction
DPO is an alignment method that does not require an auxiliary reward model, thereby improving training efficiency and stability. DPO training reduces to logistic regression on offline preference datasets, following the Bradley–Terry model bradley1952rank . Subsequent studies have proposed extensions of the DPO loss functions tang2024generalized ; wang2024beyond ; xu2025fpo . From the perspective of distribution matching, $f$-PO ji2024towards ; xu2025fpo reformulates the alignment objective as matching the policy model to the optimal policy defined by DPO, extending the loss using $f$-divergence. While this reformulation retains the target optimality, $f$-PO introduces additional complexity since it relies on a learned reward model and requires Monte Carlo estimation of partition functions. Importantly, its optimality cannot be guaranteed without incurring additional computational cost, as summarized in the third and fourth rows of Table˜1.
@@
We empirically evaluate several instances of BPO against baselines, measuring both generation fidelity (win rate with GPT-4 judge) and generation diversity. In contrast to other probabilistic loss extensions such as $f$-DPO wang2024beyond and $f$-PO xu2025fpo , which show a clear trade-off between win rate and diversity, BPO instances improve both win rate and entropy over DPO, as summarized in Figure˜1. The proposed SBA instance achieves the best trade-off between win rate and diversity among all instances. We also apply the proposed loss to Llama-3-8B-Instruct grattafiori2024llama and achieve a 55.9% length-controlled win rate against GPT-4 Turbo on AlpacaEval2 alpaca_eval . To the best of our knowledge, this is state-of-the-art performance among Llama-3-8B backbone models.
Table 1: Summary of probabilistic DPO extensions. O: optimality preservation, S: simplicity without extra training cost, G: generality for multiple objectives. Refer to Appendix˜B for baseline details.
@@ 2.1 Alignment of large language models with human preference
[…]
@@
Direct preference optimization (DPO): The alignment method DPO rafailov2023direct unifies the objectives Eqs.˜1 and 2 into a single objective without the reward model. Eq.˜2 has a closed-form solution:
@@
where $Z({\mathbf{x}})=\sum_{{\mathbf{y}}}\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})\text{exp}\big(\frac{1}{\beta}r_{\phi^{*}}({\mathbf{x}},{\mathbf{y}})\big)$ is the partition function. From Eq.˜3, DPO defines a new reward model $r_{\text{DPO}}({\mathbf{x}},{\mathbf{y}})=\beta\log\frac{\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})}+\beta\log Z({\mathbf{x}})$, which is parameterized by the policy and reference models. Substituting $r_{\text{DPO}}$ into $\mathcal{L}_{\text{reward}}(r_{\text{DPO}};p_{\text{data}})$ yields the following objective:
@@
which enables the policy model to learn directly from a preference dataset in a supervised manner. The optimal policy $\pi_{\theta^{*}}$ that minimizes $\mathcal{L}_{\text{DPO}}$ matches the optimal solution in Eq.˜3, and it satisfies:
$p_{\textrm{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}|{\mathbf{x}})=\sigma\left(\beta\log{\frac{\pi_{\theta^{*}}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\textrm{ref}}({\mathbf{y}}_{w}|{\mathbf{x}})}}-\beta\log{\frac{\pi_{\theta^{*}}({\mathbf{y}}_{l}|{\mathbf{x}})}{\pi_{\textrm{ref}}({\mathbf{y}}_{l}|{\mathbf{x}})}}\right).$ (5)
@@ 2.2 Likelihood ratio estimation under Bregman divergence
where $h$ denotes a strictly convex and twice continuously differentiable function with derivative function $h^{\prime}$, and $B_{h}$ is the pointwise Bregman divergence which measures the error of the linear approximation as illustrated in Figure˜2. Specific instances are summarized in Table˜2, showing that previous likelihood ratio estimation methods differ only by the choice of $h$. Among various instances, Basu’s power (BA) divergence basu1998robust , defined for $\lambda>-1$, smoothly interpolates between KLIEP (at $\lambda=0$) and LSIF (at $\lambda=1$). Inspired by the success of generalized likelihood ratio estimation in recent generative modeling studies lou2024discrete ; pmlr-v202-kim23i ; kim2024training , we apply this extension to generalize the DPO loss.
@@ 3 Methods
This section introduces the proposed Bregman Preference Optimization (BPO). Section˜3.1 reformulates the DPO objective as a likelihood ratio estimation problem. Section˜3.2 extends the DPO objective via Bregman divergence under the likelihood ratio estimation perspective. Section˜3.3 analyzes instances of BPO and introduces the scaled Basu’s power (SBA) divergence within this framework. Finally, Section˜3.4 discusses the applicability of BPO to other DPO variants in an orthogonal manner.
@@ 3.1 Preference optimization as likelihood ratio estimation
In contrast to prior works ji2024towards ; xu2025fpo that characterize the optimal policy using a learned reward model $r_{\phi^{*}}$ and a partition function $Z({\mathbf{x}})$ as in Eq.˜3, we aim to characterize the optimal policy without such modeling complexity.
@@
Please refer to Section˜A.1 for the proof. Eq.˜7 shows that the likelihood ratio of optimal policy $\pi_{\theta^{*}}$ can be specified solely using the reference model $\pi_{\text{ref}}$ and the preference data distribution $p_{\text{data}}$. The ratio $\frac{\pi_{\theta^{*}}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\theta^{*}}({\mathbf{y}}_{l}|{\mathbf{x}})}$ is equivalent to the concrete score meng2022concrete ; lou2024discrete (up to a constant), and the concrete score satisfies the completeness property lyu2012interpretation ; meng2022concrete . This means that the concrete score can uniquely identify the distribution $\pi_{\theta^{*}}$. Therefore, the matching $\frac{\pi_{\theta}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\theta}({\mathbf{y}}_{l}|{\mathbf{x}})}$ to $\frac{\pi_{\theta^{*}}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\theta^{*}}({\mathbf{y}}_{l}|{\mathbf{x}})}$ is sufficient for $\pi_{\theta}$ to recover $\pi_{\theta^{*}}$, providing a theoretical justification for ratio matching. To build a data-driven estimator for this matching, we rearrange Eq.˜7, as detailed in Eq.˜15 to Eq.˜19, and reformulate preference optimization as a matching problem between data ratio $R_{\text{data}}$ and model ratio $R_{\theta}$, defined as:
@@ 3.2 Bregman preference optimization via ratio matching
As discussed in Section˜3.1, preference optimization can be formulated as a ratio matching between $R_{\text{data}}$ and $R_{\theta}$. We define the loss using the Bregman ratio matching framework in Section˜2.2 as:
@@
Please see Section˜A.2 for the proof. Theorem˜2 implies that the objective $D_{h}(R_{\text{data}}||R_{\theta})$ guarantees target optimality under a valid $h$, but it is intractable since evaluating $R_{\text{data}}$ at a given point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ is infeasible. As preference optimization provides only samples from $p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}\mid{\mathbf{x}})$ without direct access to the distribution, we propose a tractable alternative inspired by implicit score matching hyvarinen2005estimation :
@@
The proof is provided in Section˜A.3. The intractable term $R_{\text{data}}$ does not appear in $\mathcal{L}^{h}_{\textrm{BPO}}$, and the statistical information of $R_{\text{data}}$ is captured entirely via samples from $p_{\textrm{data}}$. As the value of function $R_{\theta}$ at the point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ computed through the feed-forward passes of $\pi_{\theta}$ and $\pi_{\text{ref}}$, $R_{\theta}$ is appropriate to retain it inside the expectation. To analyze the learning dynamics of $\mathcal{L}^{h}_{\textrm{BPO}}(R_{\theta};p_{\textrm{data}})$, we provide the following gradient analysis:
@@
See Section˜A.4 for the proof. Eq.˜11 shows that the gradient direction at a point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ depends only on $\nabla_{\theta}R_{\theta}$, regardless of the choice of $h$. In contrast, $h$ controls the point-wise gradient magnitude $G_{h}(R_{\theta})$, which determines the relative weighting of each sample during optimization. The weighting can influence the gradient direction aggregated over a mini-batch as shown in Figure˜3(c). Proposition˜4 further explains that gradient descent updates decrease the value of $R_{\theta}$ at the observed point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})\sim p_{\text{data}}$. This is intuitive since the observed pair is sampled from the distribution in the denominator of the target ratio $R_{\text{data}}=\frac{p_{\text{data}}({\mathbf{y}}_{w}\prec{\mathbf{y}}_{l}\mid{\mathbf{x}})}{p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}\mid{\mathbf{x}})}$. Figures˜3(a) and 3(b) show that $G_{h}(R_{\theta})$ remains positive across various choices of $h$. Although Theorem˜2 shows that the theoretical optimality holds regardless of the choice of the function $h$, the gradient magnitude $G_{h}(R_{\theta})$ varies with $h$, making the choice of $h$ important due to potential sub-optimality in practical optimization.
@@ 3.3 Instances of Bregman preference optimization
Table˜2 summarizes well-known instances of Bregman divergences unified in sugiyama2012density and their corresponding BPO objectives $\mathcal{L}^{h}_{\text{BPO}}$ determined by the choice of $h$.
BPO recovers DPO as a special case: The proposed $\mathcal{L}^{h}_{\text{BPO}}$ recovers the original DPO objective $\mathcal{L}_{\text{DPO}}$ when $h$ corresponds to the logistic regression, denoted as $\mathcal{L}^{\text{LR}}_{\text{BPO}}$ (See Section˜A.5 for more details):
$=\mathbb{E}_{p_{\textrm{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}|{\mathbf{x}})}\left[-\log{\sigma\left(\beta\log{\frac{\pi_{\theta}({\mathbf{y}}_{w}|{\mathbf{x}})}{\pi_{\textrm{ref}}({\mathbf{y}}_{w}|{\mathbf{x}})}}-\beta\log{\frac{\pi_{\theta}({\mathbf{y}}_{l}|{\mathbf{x}})}{\pi_{\textrm{ref}}({\mathbf{y}}_{l}|{\mathbf{x}})}}\right)}\right]=\mathcal{L_{\textrm{DPO}}}(\pi_{\theta};\pi_{\textrm{ref}},p_{\textrm{data}}).$
KLIEP and LSIF under Basu’s power divergence (BA): The instances, $\mathcal{L}^{\text{KLIEP}}_{\text{BPO}}$ and $\mathcal{L}^{\text{LSIF}}_{\text{BPO}}$, are unified under $\mathcal{L}^{\text{BA}_{\lambda}}_{\text{BPO}}$, where the gradient magnitude is given by:
@@
as shown in Figure˜3(a) for various values of $\lambda$. While the term $(R_{\theta}^{\lambda}+R_{\theta}^{-\lambda-1})$ provides meaningful control over the optimization behavior with respect to $R_{\theta}$ via the hyperparameter $\lambda$, the coefficient $(\lambda+1)$ unnecessarily scales up the gradient magnitude without introducing any $\theta$-dependent learning signal. Compared to the gradient magnitude of DPO, given by $G_{\text{LR}}(R_{\theta})=\frac{1}{1+R_{\theta}}$, the gradient magnitude of $G_{\text{BA}_{\lambda}}(R_{\theta})$ increases significantly as $\lambda$ becomes larger. When the gradient scale varies, training requires careful adjustment of the learning rate, batch size, and optimizer. Since preference optimization is sensitive to hyperparameters, improper tuning can severely degrade performance. To mitigate this issue, we propose a simple scaled version of BA to match DPO’s gradient scale.
Scaled Basu’s power divergence (SBA): We propose $h(R)=\frac{R^{1+\lambda}-R}{s\lambda(\lambda+1)}$, which corresponds to scaling the $h$-function of BA by a factor of $s(\lambda+1)$, where $s$ is a scaling constant. This scaling directly affects the gradient magnitude as follows:
@@
$G_{\text{SBA}_{\lambda}}(R_{\theta})$ eliminates the unnecessary $\lambda$-dependent amplification, resulting in a more reasonable gradient scale in Figure˜3(b). Since preference optimization typically initializes $\pi_{\theta}$ as $\pi_{\text{ref}}$, the value of $R_{\theta}$ is 1 for all input points $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ at the start of training. By setting $s=4$, the gradient magnitude of $G_{\text{SBA}_{\lambda}}$ matches that of $G_{\text{LR}}$ at initialization ($R_{\theta}=1$). The hyperparameter $\lambda$ allows the model to control whether to prioritize updates for more confident (e.g., $R_{\theta}\ll 1$ ) or less confident (e.g., $R_{\theta}\approx 1$) samples. Recent studies wu2024beta ; kim2025spread have shown that DPO is sensitive to data quality and that applying confidence-based adjustments to individual samples can be heuristically effective. SBA controls the sensitivity to confident samples by tuning $\lambda$, while preserving theoretical optimality.